# TREACLE: Thrifty Reasoning via Context-Aware LLM and Prompt Selection

### v1

@@ 5 Experiments
We first describe the experiment setup (§5.1) and then the main results (§5.2). Specifically, we examine robustness to new LLMs and changing API prices (§5.2.1), time-varying API query latency (§5.2.2), shifts in question difficulty (§5.2.3), and different reasoning datasets (§5.2.4).
@@ 5.1 Experiment Setup
5.1.1 Datasets
We use three representative datasets for the experiments.
  • GSM8K (Cobbe et al., 2021): The Grade School Math 8K dataset contains 8.5K high quality grade school math problems created by human writers, in which 7.5K are in the training data and 1K are in the testing data. We further split the 7.5K training data into 6K training data and 1.5K validation data for calibration.
  • CSQA (Saha et al., 2018): The Complex Sequential Question Answering dataset consists of 12102 multiple choice commonsense reasoning questions encountered in daily life. The training set, validation set, and testing set contain 9741, 1221 and 1140 samples respectively.
  • LLC (Wei et al., 2022) The Last Letter Concatenation task is to concatenate the last letters of words in a name (e.g., “Amy Brown” $\rightarrow$ “yn”).
5.1.2 Data collection and training
(1) Collecting query-response pairs.
We collect query-response pairs from each dataset for different combinations of LLM, prompt, and LLM temperature. The accuracy, latency, and monetary price of the best combinations are shown in Figure 3, with full results in Table 4 in the Appendix. We selected those combinations at the Pareto frontier of accuracy and cost. Generally, larger models and more sophisticated prompts significantly improve accuracy, but with a higher financial cost.
LLMs. We used 5 different LLMs: Llama-2-7b-chat, Llama-2-13b-chat (Touvron et al., 2023), GPT-3.5-turbo, GPT-4, and GPT-4-turbo (OpenAI, 2023). These models are of varying sizes (7b, 13b, 154b and 1.76t respectively). The Llama models are open-source and run locally on our servers (one A40 GPU for Llama-2-7b and two A40 for Llama-2-13b), while the GPT models rely on commercial APIs.
Prompt types. We employ several prompting schemes to elicit the reasoning abilities of LLMs. The full prompts are shown in Appendix E. A prompt generally consists of two parts: the “content message” containing the question, and the “system message” with additional context. The full prompts are given in Appendix E.
@@ 5.1.2 Data collection and training
  • The standard few-shot prompt includes a system message (“Follow the given examples and answer the question” (Wei et al., 2022)) and the content message, which consists of few-shot examples together with the plain text prompt. It tends to improve response accuracy compared to the plain text prompt.
  • The Chain-of-Thought (CoT) few-shot prompt (Wei et al., 2022) adds some intermediate explanations to the few-shot examples.
(2) Training TREACLE.
We used Deep Q-Network (DQN) (Mnih et al., 2015) to train the reinforcement learning (RL) policy in TREACLE, consisting of a two-layer neural network. To generate diverse trajectories consisting of $(s_{t},a_{t},r_{t},s_{t+1})$, we use the collected query-response data and employ $\epsilon$-greedy exploration. For the monetary prices, we use the published per-token prices for the GPT models. Since our local Llama deployments do not have API costs, we set the Llama-2-7b’s price as $\alpha$ times Llama-2-13b’s price, and Llama-2-13b’s price as $\alpha$ times GPT-3.5-turbo’s price. $\alpha$ varies between $\frac{1}{10},\frac{1}{20}$ or $\frac{1}{50}$. Our pricing is grounded in reality and similar to actual market rates, as the offered price for Llama is approximately 15% of GPT-3.5-turbo according to current providers (together.ai, ). For the latency-accuracy tradeoff, we evaluate different trade-off parameters $\beta=[50\text{k},500\text{k},1\text{M}]$ in the cost function. We set the smallest $\beta$ to 50k because then the two terms in the cost function have similar order of magnitude. The details of the cost function values are provided in Table 5 in the Appendix. Unless stated otherwise, we set $\lambda=5$.
@@
5.1.3 Baselines
  • FrugalGPT (Chen et al., 2023). We reproduce FrugalGPT, which uses a DistilBERT model (Sanh et al., 2019) to estimate the response accuracy. If this estimate is below a threshold, the next LLM in the cascade is queried. This baseline shows how TREACLE compares to the state-of-the-art that lacks re-querying.
  • Majority Voting. For each query, we output the final response based on the majority vote from $N$ re-queries, based on (Wang et al., 2022; Yue et al., 2023). We set $N=2$ based on the best empirical results. The (LLM, prompt) combinations are progressively queried until their per-question budget runs out. This baseline allows comparison with TREACLE’s response consistency feature in the state vector.
  • Offline and online knapsack. Given the cost of LLM responses and their accuracy, we formulate a multiple choice knapsack problem where the items are the $(M+P)$ combinations, the values are the correctness probabilities, and the costs are the latency and monetary price functions. Solving this offline knapsack problem gives the optimal solution when re-queries are not allowed. We also implement an online approximation algorithm (Chakrabarty et al., 2008). These baselines show how TREACLE compares to methods with perfect knowledge of question costs and accuracy.
5.2 Results
To evaluate the performance of ${\sf\small TREACLE}$, we conduct experiments for different total budgets and cost functions. The results are presented in Figure 4. (Due to space limitations, we focus on GSM8K here; results on CSQA and LLC are similar and relegated to Section C.2.) Across different settings of $\alpha$, $\beta$, and total budget, TREACLE consistently outperforms the baselines and is close to the Offline Knapsack– an approach not feasible in practical deployments. We note that the relatively good performance of the Calibrated Cascade is due to it using the same state vector we designed for TREACLE. To delve deeper into the behavior of TREACLE, we perform additional experiments and make the following observations.
$\bullet$ Observation 1: TREACLE can adapt to different budgets and cost parameters. All the results in Figure 4, with different budgets and $\alpha,\beta$ parameters were produced after training TREACLE only once (with different parameter settings during the training). This highlights TREACLE’s adaptability to different cost function variations.
$\bullet$ Observation 2: For limited budgets, TREACLE only answers questions that are more likely to produce accurate responses. For example, for GSM8K ($\beta=50\text{k}$) in Figure 4, when the budget is only \$$\$0.05$ and insufficient for all queries, 52.7% of the questions TREACLE chooses to answer are correct. To provide context, the cheapest model (Llama-2-7b) can only answer 23.65% of questions correctly. Thus TREACLE can evaluate question difficulty and opt not to respond to those that are infeasible within the budget.
$\bullet$ Observation 3: For larger budgets, ${\sf\small TREACLE}$ intelligently chooses more powerful model-prompt combinations. This is shown in Figure 5(a), which shows the average number of times each model-prompt combination is re-queried. As the budget increases, the more powerful models (right side of x-axis) are increasingly selected). Interestingly, we observe that for budgets \$0.3 to \$10, the Llama-2-13b model is queried approximately once per question, despite its suboptimal performance. Even with these larger budgets, it’s still beneficial to check Llama before moving onto more powerful models, to see whether its responses are consistent.
@@ 5.2 Results
$\bullet$ Observation 4: Re-querying helps. We conducted an ablation study where we trained both TREACLE and the Calibrated Cascade Algorithm baseline without the ability to re-query. The results are shown in Figure 6, where the dashed line represents method variants that permits re-querying. We observed a notable decrease in accuracy when re-querying was not allowed. Methods without re-querying eventually achieved comparable accuracy with those with re-querying capability, but with significantly larger budgets.Additional ablation experiments showing that re-querying or prompt selection help are shown in Section C.3.
@@
Figure 6: Performance of ${\sf\small TREACLE}$ with and without the ability to re-query. The dashed lines represent the methods that allow re-querying. Experiments with $\alpha=1/20$.
@@ 5.2.1 Addition of new LLMs
LLM development is rapid, with better models continuously emerging, and the API prices set by providers can change at any time. TREACLE’s ability to react to such changes is thus an important practical consideration. We show that TREACLE can adapt by fine-tuning itself using few samples.
We study two types of changes to the LLMs and their prices. (1) API price adjustment: In November 2023, OpenAI released GPT-4-turbo, offering performance on par with GPT-4 but at a more affordable price. Concurrently, the price for GPT-3.5-turbo was lowered. (2) Fine-tuned open-source LLMs: Several domain-specific fine-tuned models with higher accuracy have been released. Specifically, we exchanged Llama-2 for MetaMath (Yu et al., 2023), which is fine-tuned specifically for GSM8K. For both scenarios, we partitioned the GSM8K test data into 80% validation and 20% test samples, generated new state-action trajectories from the validation set, then fine-tuned TREACLE on these new trajectories. To create a comparable baseline, we similarly fine-tuned FrugalGPT’s DistilBERT.
@@
Figure 7: Performance with new LLMs and lowered prices. Lines and dots in light (dark) colors are results with old (new) prices and LLMs.
@@
Firstly, we show the performance of TREACLE with both the API price adjustments and improved LLMs in Figure 7. The individual points on the plot illustrate the changes in the API prices for gpt-3.5-turbo. The lines show the performance of the new TREACLE with new models and prices and the old TREACLE (i.e., from previous subsections). The results shows that the new TREACLE can achieve the peak accuracy with only a \$1 budget, clearly benefiting from the new models and lowered prices. Benefits are also significant for lower budgets, where the improved TREACLE has significantly higher accuracy, because the lowest performing Llama-2 models were replaced by fine-tuned Metamaths. Finally, for a FrugalGPT that relies on a fine-tuned DistilBERT accuracy estimator, performance didn’t improve and can even degrade due to distribution shifts and overfitting.
(a) With new GPT models and prices (API price adjustment)
(b) With new Llama models (fine-tuned open-source LLMs)
@@
Figure 8: Sample complexity for different budgets.

### v2

@@ 5 Experiments
We first describe the experiment setup (§5.1) and then the main results (§5.2). Specifically, we examine robustness to new LLMs and changing API prices (§5.2.1), shifts in question difficulty (§5.2.2), and different reasoning datasets (§5.2.3).
@@ 5.1 Experiment Setup
We summarize the experiment setup, with full details in Appendix A. We use three representative datasets: GSM8K [4], which contains 8.5K high quality grade school math problems created by human writers; CSQA [11], which consists of 12102 multiple choice commonsense reasoning questions encountered in daily life; and LLC [17], where the task is to concatenate the last letters of words in a name (e.g., “Amy Brown” $\rightarrow$ “yn”). To evaluate our methods, we perform two steps.
(1) Collect query-response pairs for (LLM, prompt) combinations. We collected query-response pairs from each dataset for different combinations of LLM, prompt, and LLM temperature. We used 5 different LLMs: Llama-2-7b-chat, Llama-2-13b-chat [15], GPT-3.5-turbo, GPT-4, and GPT-4-turbo [10]. These models are of varying sizes (7b, 13b, 154b and 1.76t respectively). The Llama models are open-source and run locally on our servers, while the GPT models rely on commercial APIs. We employ several prompting schemes. A prompt generally consists of two parts: the “content message” containing the question, and the “system message” with additional context.
  • $\bullet$
    The plain text prompt submits the questions to the LLM as the content message.
  • $\bullet$
@@
  • $\bullet$
    The standard few-shot prompt includes a system message (“Follow the given examples and answer the question” [17]) and the content message, which consists of few-shot examples together with the plain text prompt.
  • $\bullet$
    The Chain-of-Thought (CoT) few-shot prompt [17] adds some intermediate explanations to the few-shot examples.
(2) Train TREACLE with the query-response pairs. We used Deep Q-Network (DQN) [8] to train the reinforcement learning (RL) policy in TREACLE, consisting of a two-layer neural network. For the monetary prices, we use the published per-token prices for the GPT models. Since our local Llama deployments do not have API costs, we set Llama-2-7b’s price as $\alpha$ times Llama-2-13b’s price, and Llama-2-13b’s price as $\alpha$ times GPT-3.5-turbo’s price. $\alpha$ varies between $\frac{1}{10},\frac{1}{20}$ or $\frac{1}{50}$. Our pricing is grounded in reality and similar to actual market rates, as the offered price for Llama is approximately 15% of GPT-3.5-turbo according to current providers [14]. For the latency-accuracy tradeoff, we evaluate different trade-off parameters $\beta=[50\text{k},500\text{k},1\text{M}]$ in the cost function.
  • $\bullet$
    FrugalGPT [3]. We reproduce FrugalGPT, which uses a DistilBERT model [12] to estimate the response accuracy. If this estimate is below a threshold, the next LLM in the cascade is queried. This baseline shows how TREACLE compares to the state-of-the-art that lacks re-querying.
  • $\bullet$
    Calibrated cascade. We build on FrugalGPT’s response accuracy estimation and develop a 2-layer neural network, which takes as input TREACLE’s state vector and outputs the estimated response accuracy. If this estimated accuracy is below a threshold, the next LLM in the cascade is queried. This baseline compares TREACLE to a modified FrugalGPT.
  • $\bullet$
    Majority Voting. For each query, we output the final response based on the majority vote from $\Omega$ re-queries, based on [16, 20]. We set $N=2$ based on the best empirical results. The (LLM, prompt) combinations are progressively queried until their per-question budget runs out. This baseline allows comparison with TREACLE’s response consistency feature in the state vector.
  • $\bullet$
    Offline and online knapsack. We formulate a multiple choice knapsack problem where the items are the $(M+P)$ combinations. We solve the offline knapsack to find the optimal solution when re-queries are not allowed, and also implement an online version [2]. These baselines show how TREACLE compares to methods with perfect knowledge of question costs and accuracy.
5.2 Results
@@
To evaluate the performance of ${\sf\small TREACLE}$, we conduct experiments for different total budgets and cost functions. The results are presented in Figure 3 (additional results on CSQA and LLC are in Section C.4.) Across different settings of $\alpha$, $\beta$, and total budget, TREACLE consistently outperforms the baselines and is close to the Offline Knapsack– an approach not feasible in practical deployments. We note that the relatively good performance of the Calibrated Cascade is due to it using the same state vector we designed for TREACLE. We make the following additional observations.
$\bullet$ Observation 1: TREACLE can adapt to different budgets and cost parameters. All the results in Figure 3, with different budgets and $\alpha,\beta$ parameters were produced after training TREACLE only once. This highlights TREACLE’s adaptability to different cost function variations.
$\bullet$ Observation 2: For limited budgets, TREACLE only answers questions that are more likely to produce accurate responses. For example, for $\beta=50\text{k}$ in Figure 3, when the budget is only \$$\$0.05$ and insufficient for all queries, 52.7% of the questions TREACLE chooses to answer are correct. For context, the cheapest model (Llama-2-7b) can only answer 23.65% of questions correctly. This suggests TREACLE can evaluate question difficulty and opt not to respond to some questions.
@@
$\bullet$ Observation 3: For larger budgets, ${\sf\small TREACLE}$ chooses more powerful (LLM, prompt) combinations. This is shown in Figure 4(a), where as the budget increases, the more powerful models (right side of x-axis) are increasingly selected. Interestingly, we observe that for budgets \$0.3 to \$10, the Llama-2-13b model is queried approximately once per question, despite its suboptimal performance. Even with these larger budgets, it’s still beneficial to query Llama before moving onto more powerful models, to see whether its responses are consistent.
Figure 5: With and without re-querying. $\alpha=\frac{1}{20}$.
$\bullet$ Observation 4: Re-querying helps. We trained both TREACLE and the Calibrated Cascade Algorithm baseline without the ability to re-query. The results are shown in Figure 5, where the dashed line represents method variants that permits re-querying. We observed a notable decrease in accuracy when re-querying was not allowed. Methods without re-querying eventually achieved comparable accuracy with those with re-querying capability, but with significantly larger budgets.
@@ 5.2.1 Addition of new LLMs
Figure 6: Performance with new LLMs and lowered prices. Lines and dots in light (dark) colors are results with old (new) prices and LLMs. $\alpha=\frac{1}{10}$.
LLM development is rapid, with better models continuously emerging, and the API prices set by providers can change at any time. TREACLE’s ability to react to such changes is thus an important practical consideration. We show that TREACLE can adapt by fine-tuning itself using few samples. We study two types of LLM updates. (1) API price adjustment: In November 2023, OpenAI released GPT-4-turbo, offering performance on par with GPT-4 but at a more affordable price. Concurrently, the price for GPT-3.5-turbo was lowered. (2) Fine-tuned open-source LLMs: Several domain-specific fine-tuned models with higher accuracy have been released. Specifically, we exchanged Llama-2 for MetaMath [19], which is fine-tuned specifically for GSM8K. For both scenarios, we partitioned the GSM8K test data into 80% validation and 20% test samples, generated new state-action trajectories from the validation set, then fine-tuned TREACLE on these new trajectories. To create a comparable baseline, we also fine-tuned FrugalGPT’s DistilBERT.
@@
Firstly, we show the performance of TREACLE with both the API price adjustments and improved LLMs in Figure 6. The individual points on the plot illustrate the changes in the API prices for gpt-3.5-turbo. The lines show the performance of the new TREACLE with new models and prices and the old TREACLE (from previous subsections). The new TREACLE can achieve the peak accuracy with only a \$1 budget, clearly benefiting from the new models and lowered prices. Benefits are also significant for lower budgets, where the improved TREACLE has significantly higher accuracy, because the lowest performing Llama-2 models were replaced by fine-tuned Metamaths. Finally, for FrugalGPT that relies on a fine-tuned DistilBERT accuracy estimator, performance didn’t improve and can even degrade due to distribution shifts and overfitting.
(a) New GPT models and prices (API price adjustment)
(b) New Llama models (fine-tuned open-source LLMs)
@@
Figure 7: Sample complexity for different budgets with new LLMs. $\alpha=\frac{1}{10}$.